A particle with charge 2 mu C is located on the x-
Answer #1

The electrostatic force is given by

F=k_e\frac{|q_1||q_2|}{d^2}

The charge in +10cm push to the left the charge at -2cm. The distance between them is d=10cm-(-2cm)=12cm= 0.12m

The charge in -10cm pull to the left the charge at -2cm. The distance between them is d=-2cm-(-10cm)=8cm= 0.08m

Since the to forces has the same direction the net force in the charge at at -2cm is

F=(8.9875\times 10^9N\cdot m^2/C^2)\frac{(2\times 10^{-6}C)(7\times 10^{-6}C)}{(0.08m)^2}+(8.9875\times 10^9N\cdot m^2/C^2)\frac{(9\times 10^{-6}C)(7\times 10^{-6}C)}{(0.12m)^2}

F=58.98N

Toward left